OpenStudy (anonymous):

At Trent's garage, 3 cases of brake fluid last 4 weeks. Trent has 5 cases. How much more brake fluid should Trent order so that he has enough for 13 weeks? Round your answer up to the nearest whole number.

OpenStudy (anonymous):

this is simple

OpenStudy (anonymous):

first find how long a singles case lasts thats is 4/3= 1.3

OpenStudy (anonymous):

now he already has 5 cases, so we should subtract that from the answer

OpenStudy (anonymous):

13/1.3 = ?

OpenStudy (anonymous):

can you figure that out?

OpenStudy (anonymous):

the equation is (13/1.33) - 5 = x

OpenStudy (anonymous):

use your calculator

OpenStudy (anonymous):

or you can use (13/4 x 3) - 5 = x

OpenStudy (anonymous):

whichever you choose

OpenStudy (anonymous):

then the answer is rounded by the hundredth place

OpenStudy (halo):

thanks

OpenStudy (anonymous):

the answer is 5 cases